OpenStudy (anonymous):

How to find the square root of 900 the fast way? ( without using the caculator)

OpenStudy (anonymous):

calculator*

OpenStudy (anonymous):

well the fast way would be to use the calculator but you would simplify it to \[\sqrt{9}\sqrt{10}\]

OpenStudy (anonymous):

maybe square root of nine is three, square root of one hundred is ten, three times ten is thirty

OpenStudy (anonymous):

i don't get it?

OpenStudy (anonymous):

sqrt900 =sqrt(9*100) =sqrt9*sqrt100 =3*10 =30

OpenStudy (anonymous):

\[\sqrt{9}\sqrt{100}\] sorry so this means 3x10 which is 30

OpenStudy (anonymous):

Yes i get it now...........